Sasha owns two investments, A and B, that have a combined total value of 31,300 dollars. Investment A is expected to pay 28,600 dollars in 7 year(s) from today and has an expected return of 16.12 percent per year. Investment B is expected to pay 35,676 in 6 years from today and has an expected return of R per year. What is R, the expected annual return for investment B? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Let the amount invested in A be x

Future value of x at 16.12% interest rate in 7 years is 28600

=> 28600 = x (1+16.12/100)7

=> x = 10046.55

Total Investment is 31300

=> Amount invested in B = 31300 - 10046.55 = 21253.45

Rate of return for investment B = R

Future value of amount invested in B after 6 years = 35676

=> 35676 = 21253.45(1+R)6

=> R = 0.09016

Hence, R = 9.02%